OpenStudy (anonymous):

Two jets leave Tulsa at 3:00 p.m. traveling in opposite directions. The northbound jet is flying at a rate of 850 km/h and the southbound jet is flying at a rate of 750 km/h. At what time will they be 4000 km apart?

OpenStudy (anonymous):

Distance = (net velocity) (time) =(850 + 750) (time) 4000 km = (850 + 750 km/h) (time)